Can you explain to me how to find five ordered pairs to equations like y=3x + 1?
Found 2 solutions by scott8148, Edwin McCravy:
Answer by scott8148(6628)

ordered pairs are solutions of the equation

pick 5 values for x and solve for the 5 corresponding y values

the pairs are written as (x value , y value)

Answer by Edwin McCravy(20086)

y = 3x + 1
Pick a number for the x-coordinate.
I will pick, say, x = -3. Now I find
the y-coordinate by substituting (-3)
for x in
y = 3x + 1
y = 3(-3) + 1
y = -9 + 1
y = -8
So one ordered pair has its x-coordinate as
-3 and its y-coordinate as -8. So this ordered
pair is (-3,-8).
---------------------
Pick another number for the x-coordinate.
I will pick, say, x = 2. Now I find
the y-coordinate by substituting (2)
for x in
y = 3x + 1
y = 3(2) + 1
y = 6 + 1
y = 7
So another ordered pair has its x-coordinate as
2 and its y-coordinate as 7. So this ordered
pair is (2,7).
---------------------
Pick another number for the x-coordinate.
I will pick, say, x = 0. Now I find
the y-coordinate by substituting (0)
for x in
y = 3x + 1
y = 3(0) + 1
y = 0 + 1
y = 1
So another ordered pair has its x-coordinate as
0 and its y-coordinate as 1. So this ordered
pair is (0,1).
---------------------
Pick another number for the x-coordinate.
I will pick, say, x =-2. Now I find
the y-coordinate by substituting (-2)
for x in
y = 3x + 1
y = 3(-2) + 1
y = -6 + 1
y = -5
So another ordered pair has its x-coordinate as
-2 and its y-coordinate as -5. So this ordered
pair is (-2,-5).
---------------------
Pick another number for the x-coordinate.
I will pick, say, x = 1. Now I find
the y-coordinate by substituting (1)
for x in
y = 3x + 1
y = 3(1) + 1
y = 3 + 1
y = 4
So another ordered pair has its x-coordinate as
1 and its y-coordinate as 4. So this ordered
pair is (1,4).
---------------------
So 5 ordered pairs are (-3,-8), (2,7), (0,1), (-2,-5), and (1,4).
If you plot those on a set of axes, you will get this:

and you will notice that you can take a ruler and
draw a line and it will go through all of them, just
like magic!




Edwin